# Physics-Guided Multi-Fidelity DeepONet for Data-Efficient Flow Field Prediction

### v1

@@ 4.2 Effects of Spatial Point Sampling and Application of Automatic Mixed Precision
The results reveal important insights about DeepONet’s training efficiency through the combination of sampling strategies and AMP training. First, examining the impact of spatial point sampling, we observe that smaller sampling sizes can maintain prediction accuracy while significantly reducing computational costs. With AMP enabled, the 16×16 sampling strategy achieves comparable accuracy (test error: 17.419) to the baseline that uses all available points (test error: 17.079), while dramatically reducing the training time from 8.672 to 0.350 hours—a 96% reduction. Similar trends are observed in cases without AMP, where the 16×16 sampling rather achieves a test error of 16.876 compared to 18.001 for full sampling due to its effective sampling nature compared to naive all-point-sampling approach.
@@
The comparison between AMP and non-AMP training reveals interesting trade-offs in computational efficiency and prediction accuracy across different sampling sizes. While non-AMP training with 16×16 sampling achieves slightly better accuracy (16.876 versus 17.419 with AMP) and requires only marginally longer training time (0.377 versus 0.350 hours), we opt for the AMP configuration with 16×16 sampling in subsequent experiments. This choice is motivated by two practical considerations: (1) our comprehensive study involves extensive experiments, each repeated multiple times to account for training stochasticity—in this context, even small improvements in computational efficiency become significant when scaled across numerous training runs; and (2) AMP’s computational benefits scale dramatically with sampling density—for the full 128×128 grid, AMP reduces training time by 36.5% compared to non-AMP training. This scalable efficiency gain underscores the importance of investigating AMP in DeepONet applications, as many practical implementations may require substantially denser spatial sampling than explored here. For these reasons—both to enhance efficiency across our numerous experimental runs and to leverage the scalable computational benefits that become increasingly significant with larger sampling sizes—we employ AMP in all subsequent experiments.
@@
In Figure 8, we can observe the effects of different sampling resolutions on flow field prediction quality given AMP. The comparison demonstrates that even with the coarsest sampling resolution (16×16), the DeepONet successfully captures the essential vorticity patterns and intensity distributions present in the flow field at $t=0.25s$. As expected, the visual differences between sampling resolutions are minimal, with all configurations reproducing the key flow structures observed in the ground truth. This visual evidence supports our quantitative findings in Table 2, which indicate that the 16×16 sampling strategy achieves comparable prediction accuracy (MSE of 17.419) to the full 128×128 sampling (MSE of 17.079), while dramatically reducing the training time from 8.672 to 0.350 hours. The consistent visual quality across different sampling resolutions highlights DeepONet’s point-based prediction flexibility, which allows for efficient data utilization without substantial degradation in prediction performance.
@@
Figure 8: To visually investigate the effects of number of sampled points, 4 cases with AMP in Table 2 are visualized: snapshots at $t=0.25s$. The comparison demonstrates that even with significantly reduced sampling (16×16), the model captures essential flow patterns similar to those observed in higher resolution samplings (32×32, 64×64) and the full dataset (128×128).
@@
Note: The findings in this section highlight a fundamental advantage of DeepONet over grid-based architectures like CNNs or GNNs: its point-based prediction capability enables remarkable flexibility in data sampling, free from the constraints of fixed grid structures that limit traditional architectures and often require storing complete grid information in memory. This sampling flexibility, when combined with computational optimization techniques like AMP training, provides an effective strategy for enhancing DeepONet’s training efficiency while significantly reducing memory requirements compared to grid-based methods. Such adaptability is particularly valuable for real-world applications where computational resources may be limited, allowing users to strategically balance accuracy, memory usage, and computational cost through appropriate sampling strategies.
@@ 4.3 Effects of Positional Encoding
Building upon the optimal merge network configuration and sampling strategy identified in previous sections, merge Type 1 with 16×16 sampling with AMP, we investigate the impact of positional encoding (PE) on DeepONet’s prediction accuracy. PE transforms input coordinates using sinusoidal functions, potentially mitigating the spectral bias inherent in neural networks (Section 2.1.2). We systematically explore two distinct PE strategies: (1) PE($\mathbf{t}$): applying PE only to the temporal coordinate ($t$), and (2) PE($\mathbf{x,y,t}$): applying PE to all trunk network inputs ($x$, $y$, and $t$). For each strategy, we evaluate combinations of two key hyperparameters: the scale factor $\sigma$ controlling the frequency range, and the mapping size $m$ determining the dimension of encoded features. All other training settings remain consistent with Section 4.1.
@@
The comparison of different PE strategies reveals several important insights about DeepONet’s behavior with coordinate encoding (Table 3). First, it’s noteworthy that all models demonstrate similar training times (approximately 0.35-0.37 hours with NVIDIA A6000 GPU), indicating that the application of PE does not meaningfully impact computational efficiency. This consistent training overhead across all configurations makes PE an attractive option for potential performance improvement without computational penalties. The results show a clear distinction between temporal-only and full spatio-temporal encoding strategies. PE($\mathbf{t}$) improves model performance, with the best configuration ($\sigma=10$, $m=32$) achieving a test error of 16.100 compared to 17.419 without PE—a 7.57% improvement. In contrast, PE($\mathbf{x,y,t}$) leads to deteriorated performance across all configurations, with test errors increasing significantly as the scale factor increases. This deterioration becomes particularly severe with higher scale factors, culminating in training instability at $\sigma=10$ and $m=64$ where four out of five training runs diverged, indicating that high-frequency embeddings of spatial coordinates can destabilize the training process in our case.
@@
The impact of hyperparameters reveals important considerations for PE implementation. The scale factor $\sigma$, which controls the frequency range of the encoding, shows strong influence on model performance. In PE($\mathbf{t}$), higher $\sigma$ values generally lead to better performance, suggesting that temporal dynamics in our flow field benefit from higher-frequency encodings. However, this relationship between performance and $\sigma$ is highly dependent on the underlying temporal characteristics of the flow field, emphasizing the importance of careful hyperparameter tuning for specific applications. Furthermore, the mapping size $m$ demonstrates a non-monotonic relationship with model performance. For instance, in PE($\mathbf{t}$), increasing $m$ from 32 to 64 consistently degrades performance across all scale factors, indicating that larger mapping sizes do not necessarily translate to better predictions. This suggests an optimal intermediate dimension for the encoded features that balances expressiveness with model complexity, with $m=32$ proving to be the optimal value in our case.
@@
In conclusion, our results demonstrate that positional encoding, when properly configured and applied selectively to temporal coordinates, can enhance DeepONet’s prediction accuracy without compromising computational efficiency. The 7.57% improvement achieved with PE($\mathbf{t}$) requires no additional architectural complexity or significant implementation effort, making it a valuable addition to DeepONet implementations for spatio-temporal prediction tasks. However, practitioners should carefully consider both the encoding strategy and hyperparameter selection, since inappropriate choices—particularly in spatial coordinate encoding—can significantly impair model performance.
@@ 5.1 Effects of Different Transfer Learning Approaches with Different Low-Fidelity Dataset
    Lu et al. [18] (conventional MF-DeepONet): while the original MF-DeepONet approach proposed by Lu et al. [18] combines both residual learning and input augmentation within a coupled two-DeepONet framework, we strategically implement only the input augmentation aspect. We discard the coupled use of two DeepONet models and residual learning component due to their significant limitation: they fundamentally require identical query points across fidelity levels, which severely restricts the flexible usage of multi-fidelity datasets. Therefore, its implementation includes the sequential architecture where LF and HF DeepONets are trained separately as our proposed MF-DeepONet framework, with the LF model’s output directly feeding into the HF model’s trunk net—input augmentation aspect proposed by Lu et al. [18]. This architecture still necessitates that both networks execute in sequence during all training and inference operations, requiring much more computational time than our proposed framework.
@@
    Proposed: as proposed in Section 2.2.1, our MF-DeepONet employs fine-tuning where the pre-trained LF DeepONet’s branch and trunk networks and frozen, while only the merge network parameters are updated during HF training. This approach preserves valuable LF representations while adapting specifically to HF features.
@@
    Full-tuning: all network components (branch, trunk, and merge networks) are simultaneously trainable, potentially compromising the established low-fidelity knowledge while predominantly focusing on high-fidelity feature acquisition.
@@
    Linear probing: takes a highly restrictive approach by freezing all layers except the final layer of the merge network.
We systematically evaluate these approaches across different LF data configurations, varying both spatial resolution (16×16, 32×32, 64×64) and dataset size (50, 100, 200, 300 samples)—note that HF dataset size is fixed as 100. For fair comparison with the single-fidelity DeepONet which uses 800 training epochs, both LF and HF training phases in the multi-fidelity approaches are set to 400 epochs each, ensuring the total number of training epochs remains constant across all experiments in this paper.
The results, presented in Table 4 and also visualized as Figure 9 for easier comparison of MSE and training hour, reveal several significant insights about MF-DeepONet performance. First, our proposed fine-tuning approach consistently outperforms all other methods, including the input augmentation implementation of MF-DeepONet from Lu et al. [18]. The conventional MF-DeepONet approach (Lu et al. [18]) shows limited effectiveness, with MSE values ranging from 25.407 to 26.388 for 16×16 resolution data, significantly underperforming even single-fidelity training (MSE 16.100). While these inferior results may partly stem from our necessary omission of the residual learning component from the original Lu et al. [18] framework, this design choice was deliberate and essential for our study’s objectives. The residual learning approach, while potentially beneficial in certain contexts, fundamentally restricts DeepONet’s inherent data flexibility by requiring identical sampling points across fidelity levels—a constraint that severely limits practical applicability in real engineering scenarios where sampling locations often differ between low and high-fidelity simulations. Even at higher resolutions (64×64 LF data with 300 samples), this approach achieves an MSE of only 15.940, which is 76% higher than our fine-tuning method’s 9.057 MSE. Additionally, the sequential architecture of Lu et al. [18]’s approach—requiring both LF and HF DeepONet networks to be executed in sequence during inference—leads to consistently higher computational costs, with training times approximately 25-35% longer than our proposed method across all configurations.
@@
Table 4: Comparison of different MF-DeepONet architectures. The values highlighted in gray serve as baseline performance metrics for the analysis of time-derivative guided sampling efficiency in Section 6.
@@
Single-fidelity 0.368 16.100
@@
Second, fine-tuning consistently outperforms both full-tuning and linear probing across all configurations, demonstrating the effectiveness of our strategy to preserve learned low-fidelity latent representations by branch/trunk networks while adapting only the merge network to high-fidelity features. Particularly noteworthy is that even with the coarsest LF resolution (16×16), fine-tuning achieves better performance compared to single-fidelity training (MSE of 15.238 vs 16.100) when provided with sufficient LF data (200 samples). This improvement is remarkable given that the 16×16 resolution data exhibits significantly different flow physics compared to higher resolutions, as evidenced in Figure 6. While full-tuning provides more flexibility through complete network adaptation, this additional freedom may not be optimal, as it can lead to less efficient use of valuable low-fidelity feature representations learned during pre-training. The superior performance of fine-tuning suggests that the branch and trunk networks effectively capture essential parameter and spatio-temporal representations during low-fidelity training, which can be successfully leveraged for high-fidelity predictions through careful adaptation of the merge network alone.
@@
In terms of LF dataset, as expected, increasing the resolution of LF data from 16×16 to 64×64 leads to consistent improvements in prediction accuracy across all transfer learning approaches. With 300 LF samples at 64×64 resolution, fine-tuning achieves an MSE of 9.057, representing a 43.7% improvement over single-fidelity training (where MSE is 16.001). Figure 10 presents a comparative visualization of vorticity field predictions at two different time steps ($t=0.5s$ and $t=0.75s$). Our proposed MF-DeepONet with fine-tuning demonstrates notably superior prediction accuracy compared to the single-fidelity approach, particularly in regions highlighted by white dotted boxes. Within these regions, the single-fidelity model exhibits significant deviations from the ground truth, failing to capture the correct vorticity patterns and flow structures. In contrast, our proposed MF-DeepONet successfully reproduces the complex flow features, maintaining fidelity to the true flow dynamics across both temporal snapshots.
@@
Figure 10: Visual comparison of the predicted vorticity field at $t=0.5s$ and $t=0.75s$ between single-fidelity DeepONet and multi-fidelity DeepONet with fine-tuning (with LF data resolution of 64×64 and data size of 300) within test scenario.
Note: The results in Table 4 reveal important insights about knowledge transfer between fidelity levels. Linear probing consistently performs poorly across all configurations, with MSE values substantially higher than both fine-tuning and single-fidelity approaches. This underperformance is scientifically significant—it confirms the substantial differences between low-fidelity and high-fidelity flow physics, as the frozen representations from low-fidelity models inadequately capture high-fidelity features without adaptation. Despite these fundamental differences, our fine-tuning approach demonstrates remarkable effectiveness—even when using coarse 16×16 resolution low-fidelity data with markedly different physical characteristics from ground truth physics, the model still outperforms training on HF data alone. Specifically, fine-tuning with 16×16 resolution low-fidelity data achieves MSE values of 15.238 (with 200 samples) and 15.096 (with 300 samples), representing improvements of 5.4% and 6.2% respectively over the single-fidelity baseline (MSE: 16.100). This confirms that our selective adaptation approach effectively leverages useful information from low-fidelity simulations, even when they exhibit partially different physical characteristics compared to their high-fidelity counterparts.

### v2

@@ 4.2 Effects of Spatial Point Subsampling and Application of Automatic Mixed Precision
The comparison between AMP and non-AMP training reveals interesting trade-offs in computational efficiency and prediction accuracy across different sampling sizes. While AMP-OFF training with 16×16 sampling achieves slightly better accuracy (16.876 versus 17.419 with AMP-ON) and requires slightly longer training time (0.377 versus 0.350 hours), we opt for the AMP-ON configuration with 16×16 sampling in subsequent experiments. This choice is motivated by two practical considerations: (1) our comprehensive study involves numerous repeated experiments to account for training variability, where even small reductions in training time yield substantial cumulative savings; and (2) the computational advantage of AMP scales with sampling resolution—at full 128×128 sampling, AMP reduces training time by 36.5%. This scalable efficiency demonstrates the practical value of AMP and underscores the importance of systematically exploring its effectiveness in various DeepONet settings to be investigated in later sections. Accordingly, the AMP-ON configuration with 16×16 sampling in Table 2 is judged to represent the reasonable compromise between accuracy and computational cost in this study.
@@
In Figure 8, we can observe the effects of different subsampling resolutions on flow field prediction quality given AMP. The comparison demonstrates that even with the coarsest sampling resolution (16×16), the DeepONet successfully captures the essential vorticity patterns and intensity distributions present in the flow field at $t=0.25s$. As expected, the visual differences between sampling resolutions are minimal, with all configurations reproducing the key flow structures observed in the ground truth. This visual evidence supports our quantitative findings in Table 2, which indicate that the 16×16 sampling strategy achieves comparable prediction accuracy (MSE of 17.419) to the full 128×128 sampling (MSE of 17.079), while dramatically reducing the training time from 8.672 to 0.350 hours. The consistent visual quality across different sampling resolutions highlights DeepONet’s point-based prediction flexibility, which allows for efficient data utilization without substantial degradation in prediction performance.
@@
Figure 8: To visually investigate the effects of number of subsampled points, 4 cases with AMP in Table 2 are visualized: snapshots at $t=0.25s$. The comparison demonstrates that even with significantly reduced sampling (16×16), the model captures essential flow patterns similar to those observed in higher resolution samplings (32×32, 64×64) and the full dataset (128×128).
@@
The findings in this section highlight a fundamental advantage of DeepONet over grid-based architectures like CNNs or GNNs: its point-based prediction capability enables remarkable flexibility in data sampling, free from the constraints of fixed grid structures that limit traditional architectures and often require storing complete grid information in memory. This sampling flexibility, when combined with computational optimization techniques like AMP training, provides an effective strategy for enhancing DeepONet’s training efficiency while significantly reducing memory requirements compared to grid-based methods.
@@ 4.3 Effects of Positional Encoding
Building upon the optimal merge network configuration and sampling strategy identified in previous sections (merge Type 1 with 16×16 subsampling with AMP), we investigate the impact of positional encoding (PE) on DeepONet’s prediction accuracy for spatio-temporal flow fields. While PE is a highly effective and well-established technique for coordinate-based networks in computer vision [43, 33, 44, 57, 3, 6], its application has predominantly focused on spatial coordinates. Consequently, clear guidelines for their optimal application to complex spatio-temporal problems within operator learning frameworks like DeepONet are less established. A key motivation for this part of our study, therefore, is to systematically explore its potential effectiveness and specific behaviors when applied to spatio-temporal prediction context. This investigation aims to provide valuable insights and practical guidelines for DeepONet users working at the intersection of these domains (e.g., CFD and computer science). To this end, we systematically explore three distinct PE strategies: (1) PE($t$): applying PE only to the temporal coordinate ($t$); (2) PE($x,y$): applying PE only to the spatial coordinates ($x,y$); and (3) PE($x,y,t$): applying PE to all spatio-temporal trunk network inputs ($x,y$, and $t$). For each strategy, we evaluate combinations of two key hyperparameters: the scale factor $\sigma$ controlling the frequency range, and the mapping size $m$ determining the dimension of encoded features. Besides $\sigma$ and $m$, all other training settings remain consistent with previous settings. Though sigma values are explored in a greedy manner in this paper, please note that the selection of optimal $\sigma$ can be aided by the prior knowledge on the dataset: for example, if dominant temporal periods in the flow dynamics are known, $\sigma$ could be tuned to these frequencies to better capture temporal patterns. However, in more general cases where such prior knowledge is unavailable, a preliminary parametric study, analogous to our investigation summarized in Table 3, is advisable.
@@
The comparison of different PE strategies in Table 3 reveals several important insights about DeepONet’s behavior with coordinate encoding. First, it’s noteworthy that all models demonstrate broadly similar training times (approximately 0.35-0.46 hours with NVIDIA A6000 GPU), indicating that the application of PE does not meaningfully impact computational efficiency. However, the results show a clear distinction between encoding strategies. The temporal-only strategy, PE($t$), improves model performance, with the best configuration ($\sigma=10$, $m=32$) achieving a test error of 16.100 compared to 17.419 without PE—a 7.57% improvement. In contrast, applying PE to the spatial coordinates proved to be detrimental in our tests. The strategy of encoding all spatio-temporal inputs, PE($x,y,t$), led to deteriorated performance across all configurations. Also, experiments focusing only on spatial encoding, PE($x,y$), confirm this finding, as this strategy also consistently yielded higher errors than the baseline without PE (with a minimum test error of 20.658). This deterioration becomes particularly severe with higher scale factors for any strategy involving spatial encoding, culminating in training instability for PE($x,y,t$) at $\sigma=10$ and $m=64$, where four out of five training runs diverged.
@@
The impact of hyperparameters reveals important considerations for PE implementation. The scale factor $\sigma$, which controls the frequency range of the encoding, shows strong influence on model performance. In PE($t$), higher $\sigma$ values generally lead to better performance, suggesting that temporal dynamics in our flow field benefit from higher-frequency encodings. Conversely, for strategies involving spatial coordinates (PE($x,y$) and PE($x,y,t$)), higher $\sigma$ values dramatically worsened performance, suggesting that high-frequency spatial embeddings destabilize the training process for this problem. This relationship between performance and $\sigma$ is highly dependent on the underlying characteristics of the spatio-temporal field, emphasizing the importance of careful hyperparameter tuning for specific applications. Furthermore, the mapping size $m$ demonstrates a non-monotonic relationship with model performance. For instance, in PE($t$), increasing $m$ from 32 to 64 consistently degrades performance across all scale factors, indicating that larger mapping sizes do not necessarily translate to better predictions. This suggests an optimal intermediate dimension for the encoded features that balances expressiveness with model complexity, with $m=32$ proving to be the optimal value in our case.
@@
In conclusion, our results demonstrate that positional encoding, when properly configured and applied selectively to temporal coordinates, can enhance DeepONet’s prediction accuracy without compromising computational efficiency. The 7.57% improvement achieved with PE($t$) is a valuable addition for spatio-temporal prediction tasks, requiring no significant architectural or implementation overhead. However, the more critical insight from our systematic study is the finding that only temporal encoding is beneficial, whereas spatial encoding is detrimental for this problem. This stands in contrast to many computer vision applications where PE is successfully applied to spatial coordinates. Therefore, our investigation provides a valuable and non-obvious guideline: for spatio-temporal problems, one must carefully consider which coordinates to encode, as an indiscriminate application of PE can significantly impair model performance.
@@ 5.1 Effects of Different Transfer Learning Approaches with Different Low-Fidelity Dataset
    Lu et al. [29] (conventional MF-DeepONet): while the original MF-DeepONet approach proposed by Lu et al. [29] combines both residual learning and input augmentation within a coupled two-DeepONet framework, we strategically implement only the input augmentation aspect. We discard the coupled use of two DeepONet models and residual learning component due to their significant limitation: they fundamentally require identical query points across fidelity levels, which severely restricts the flexible usage of multi-fidelity datasets. Therefore, its implementation uses a sequential training architecture where, following the input augmentation strategy from Lu et al. [29], the output of the pre-trained LF model is used as an additional input for the HF model’s trunk network. This architecture still necessitates that both networks execute in sequence during all training and inference operations, requiring much more computational time than our proposed framework.
@@
    Proposed: as proposed in Section 2.2.1, our MF-DeepONet employs fine-tuning where the pre-trained LF DeepONet’s branch and trunk networks and frozen, while only the merge network parameters are updated during HF training. This approach preserves valuable LF representations while adapting specifically to HF features. This architecture only requires HF DeepONet during inference stage.
@@
    Full-tuning: all network components (branch, trunk, and merge networks) are simultaneously trainable, potentially compromising the established low-fidelity knowledge while predominantly focusing on high-fidelity feature acquisition. This architecture only requires HF DeepONet during inference stage.
@@
    Linear probing: takes a highly restrictive approach by freezing all layers except the final layer of the merge network. This architecture only requires HF DeepONet during inference stage.
We systematically evaluate these approaches across different LF data configurations, varying both spatial resolution (16×16, 32×32, 64×64) and dataset size (50, 100, 200, 300 samples)—note that HF dataset size is fixed as 100. For fair comparison with the single-fidelity DeepONet which uses 800 training epochs, both LF and HF training phases in the multi-fidelity approaches are set to 400 epochs each, ensuring the total number of training epochs remains constant.
The results, presented in Table 4 and also visualized as Figure 9 for easier comparison of MSE and training hour, reveal several insights about MF-DeepONet performance. First, to properly contextualize the benefits of a multi-fidelity approach, we establish the performance of single-fidelity models. Notably, training exclusively on low-fidelity data—even with a sufficient dataset size—proves insufficient for achieving competitive performance. For instance, single-fidelity models trained on 300 samples of 16×16, 32×32, and 64×64 data achieve MSE values of 70.154, 32.669, and 20.223, respectively. While the single-fidelity model trained on 100 high-fidelity samples performs better (MSE 16.100), all these single-fidelity cases are substantially outperformed by our best multi-fidelity model (MSE 9.057). This demonstrates that simply increasing the quantity of low-fidelity data cannot compensate for the fundamental need for high-fidelity information, highlighting the necessity of a multi-fidelity approach.
@@
Table 4: Comparison of different MF-DeepONet architectures. Single-fidelity results are shown for reference (e.g., “Single-fidelity (16×16, 300)” indicates 300 training samples with 16×16 resolution).
@@
Single-fidelity (16×16, 300) 0.476 70.154
Single-fidelity (32×32, 300) 1.257 32.669
Single-fidelity (64×64, 300) 1.272 20.223
Single-fidelity (128×128, 100) 0.368 16.100
@@
With these baselines established, we proceed to compare the multi-fidelity methods. Our proposed fine-tuning approach consistently outperforms all other methods, including the input augmentation implementation of the conventional MF-DeepONet (Lu et al. [29]). It shows limited effectiveness, with MSE values ranging from 25.407 to 26.388 for 16×16 resolution data, significantly underperforming even single-fidelity training with high-resolution data (MSE 16.100). Even at higher resolutions (64×64 LF data with 300 samples), this approach achieves an MSE of only 15.940, which is 76% higher than our fine-tuning method’s 9.057 MSE. While these inferior results from conventional MF-DeepONet may partly stem from our necessary omission of the residual learning component from the original Lu et al. [29] framework, this design choice was deliberate and essential for our study’s objectives. The residual learning approach, while potentially beneficial in certain contexts, fundamentally restricts DeepONet’s inherent data flexibility by requiring identical sampling points across fidelity levels—a constraint that severely limits practical applicability in real engineering scenarios where sampling locations often differ between low and high-fidelity simulations (indeed, input function parameters, $var_{1}$ and $var_{2}$, are not nested across fidelity levels in this study). Additionally, the sequential architecture of Lu et al. [29]’s approach—requiring both LF and HF DeepONet networks to be executed in sequence during inference—leads to consistently higher computational costs, with training times approximately 25-35% longer than our proposed method across all configurations.
@@
Second, fine-tuning consistently outperforms both full-tuning and linear probing across all configurations, demonstrating the effectiveness of our strategy to preserve learned low-fidelity latent representations by branch/trunk networks while adapting only the merge network to high-fidelity features. Particularly noteworthy is that even with the coarsest LF resolution (16×16), fine-tuning achieves better performance compared to single-high-fidelity training (MSE of 15.238 vs 16.100) when provided with sufficient LF data (200 samples). This improvement is unexpected given that the 16×16 resolution data exhibits remarkably different flow physics compared to higher resolutions, as evidenced in Figure 6. While full-tuning provides more flexibility through complete network adaptation, this additional freedom may not be optimal, as it can lead to less efficient use of valuable low-fidelity feature representations learned during pre-training. The superior performance of fine-tuning suggests that the branch and trunk networks effectively capture essential input function parameter and spatio-temporal representations during low-fidelity training, which can be successfully leveraged for high-fidelity predictions through careful adaptation of the merge network alone.
@@
In terms of LF dataset, as expected, increasing the resolution of LF data from 16×16 to 64×64 leads to consistent improvements in prediction accuracy across all transfer learning approaches. With 300 LF samples at 64×64 resolution, fine-tuning achieves an MSE of 9.057, representing a 43.7% improvement over single-high-fidelity training (where MSE is 16.001). Figure 10 presents their comparative visualization of vorticity field predictions at two different time steps ($t=0.5s$ and $t=0.75s$). Our proposed MF-DeepONet with fine-tuning demonstrates notably superior prediction accuracy compared to the single-fidelity approach, particularly in regions highlighted by white dotted boxes. Within these regions, the single-fidelity model exhibits significant deviations from the ground truth, failing to capture the correct vorticity patterns and flow structures. In contrast, our proposed MF-DeepONet successfully reproduces the complex flow features, maintaining fidelity to the true flow dynamics across both temporal snapshots.
@@
Figure 10: Visual comparison of the predicted vorticity field at $t=0.5s$ and $t=0.75s$ between single-fidelity (128×128 resolution) DeepONet and multi-fidelity DeepONet with fine-tuning (with LF data resolution of 64×64 and data size of 300) within test scenario.
Note on how different physics across fidelities affects multi-fidelity performance: The results in Table 4 reveal important insights about knowledge transfer between fidelity levels. Linear probing consistently performs poorly across all configurations, with MSE values substantially higher than both fine-tuning and single-fidelity approaches. This underperformance is scientifically significant—it confirms the substantial differences between low-fidelity and high-fidelity flow physics, as the frozen representations from low-fidelity models inadequately capture high-fidelity features without adaptation. Despite these fundamental differences, our fine-tuning approach demonstrates remarkable effectiveness—even when using coarse 16×16 resolution low-fidelity data with markedly different physical characteristics from ground truth physics, the model still outperforms training on HF data alone. Specifically, fine-tuning with 16×16 resolution low-fidelity data achieves MSE values of 15.238 (with 200 samples) and 15.096 (with 300 samples), representing improvements of 5.4% and 6.2% respectively over the single-fidelity baseline with 128×128 resolution (MSE: 16.100). This confirms that our selective adaptation approach effectively leverages useful information from low-fidelity simulations, even when they exhibit partially different physical characteristics compared to their high-fidelity counterparts. These results also highlight that while MF performance improves when the LF data better approximates HF physics (e.g., LF with 64×64 resolution), the proposed framework can still extract meaningful information from LF data of lower fidelity (e.g., LF with 16×16 resolution), provided the dataset is sufficiently large. This suggests that implicit correlations between LF and HF domains can still be harnessed through data-driven adaptation. Nevertheless, we still acknowledge that the effectiveness of the MF framework is sensitive to the overall quality of the LF data (Section 2.2.3), and this dependency should be carefully considered when applying the framework to problems with unreliable or weakly correlated LF sources.
Note on computational efficiency of proposed DeepONet over HF simulations: The total training time for all explored models of our proposed MF-DeepONet framework reported in Table 4 is approximately 3.958 hours, with each model averaging around 0.33 hours of training time. Considering that generation of a single CFD simulation for one HF sample (128×128 resolution) takes about 0.1 hours, the computational effort to train one MF-DeepONet model is less than generating only four new HF simulations. This comparison highlights that the one-time computational cost of training our surrogate model represents an attractive alternative to generating additional HF simulation data, since the trained DeepONet model can subsequently produce flow field predictions virtually instantaneously, enabling rapid evaluations for downstream tasks such as design optimization, flow control, or uncertainty quantification. This inference efficiency can be especially beneficial in practical scenarios where repeated evaluations are necessary and HF simulations would otherwise become a computational bottleneck.